the perimeter and area of rectangle are 22cm are 30cm square respectively. find the length and breath of the rectangle
Answer by ankor@dixie-net.com(22740)

the perimeter and area of rectangle are 22cm are 30cm square respectively.
find the length and breath of the rectangle
:
Perimeter
2L + 2W = 22
Simplify divide be 2
L + W = 11
L = (11-W)
:
Area
L * W = 30
Replace L with (11-W)
W(11-W) = 30
-W^2 + 11W = 30
-W^2 + 11W - 30 = 0
W^2 - 11W + 30 = 0 (multiplied by -1)
Factors to
(W-5)(W-6)
Two solutions
W = 5, then L = 6
W = 6, then L = 5